OpenStudy (anonymous):

Solve the system x - y = 3 and 3x - 8y = 9. What is the value of y? A. 3 B. 0 C. 6 D. -2

OpenStudy (goformit100):

B. 0

OpenStudy (anonymous):

thanks (:

OpenStudy (callisto):

x - y = 3 - (1) 3x - 8y = 9 - (2) (1) x 3 - (2) 3(x - y) - (3x - 8y) = 3x3 -9 -3y + 8y = 0 5y = 0 y = 0

OpenStudy (goformit100):

yes i was correct

OpenStudy (anonymous):

thanks guys (: